two angels are complementary. the measure of the smaller angel is 23 degrees less than the measure of the larger angel. find the measures of the angels.
Answer by jonvaliente(64)

Complimentary angles are angles whose sum is 90.
So if x=larger angle
x-23=smaller angle
and:
x + (x-23)=90
2x-23=90
2x=90+23
2x=113
x=56.5 degrees (measure of larger angle)
smaller angle=x-23=56.5-23=33.5 degrees